OpenStudy (anonymous):

cos 2θ - cos^2 θ = 0

OpenStudy (anonymous):

i cant understand your question

OpenStudy (anonymous):

use a double or half angle formula to solve the equation in the interval [0, 2pi]

OpenStudy (anonymous):

hwyne can you plz solve this

OpenStudy (anonymous):

a hollow metal sphere has an internal radius of 20 cm and an external radius of 30cm. given that the density of the metal is 7.8g/cm3 , find the mass of the sphere, your answer in kg

OpenStudy (zzr0ck3r):

\(\cos(2\theta)=\cos^2(x)-sin^2(x)\) so you have \(\cos^2(x)-\sin^2(x)-cos^2(x) = 0\\\implies -\sin^2(x) = 0\\\implies \sin(x) = 0\\\implies x=\pi*k \ \ \ \text{ where } \ \ \ k\in \mathbb{Z}\)